4. Translate 10*2to exactly 8-digit hexadecimal number. Answers: If the cost for all RAM chips to utilize the maximum memory space of the MIPS architecture is SI dolloars, how much will it cost, approximately, to implement the maximum memory space of the a 64-bit computer architecture Answer: Approximately 6 billions Filli an integer. No fraction.) 6. Fill in the missing pieces of the MIPS assembly codes that translate the following C codes int A[1001. BI 1001; B12] = h + A[41; Also, add an equivalent C code as a comment for each MIPS instruction. #Assume h is stored in $52. base addresses of A and B are in Ss6 and Ss7, respectively data A: B: space space 400 400 text lw St0. 16(So) #equivalent C code: add St0. Ss2. St0#equivalent C code: sw (Fill in the missing pieces of codes above) # equivalent C code, - - - 7. Provide a single MIPS instruction that may be used to implement the following pseudo move St0, St Answer:(Hints: consider instruction add) 8. Provide the binary representation of instruction for the following MIPS instruction sub Sv1,St1,SvO The partial solution is 0000 0000????????????0010 0010-Fill in the 16 missing binary bits in the middle. Answer OLL2.012 0221 1eo2 Sub- Description: Subtracts two registers and stores the result in a register Operation: Sd-Ss-St; advancepe (4); Syntax: sub Sd, Ss, St Encoding: 0000 00ss ssst tttt dddd d000 0010 0010
Answer #1

4)

hexadecimal = 0x2800000000

6)

lw $t0, 16($s6) # equilavent C code :- A[4]

Explanation :- A[4] = A starting address + 4 int offset = $s6 + 4 * 4 = $s6 + 16

add $t0, $s2, $t0 #equilavent C code :- h + A[4]

Explanation :- A[4] is stored in $t0 ; h is stored in $s2 .Therefore, $s2 + $t0 = h + A[4]

sw $t0, 8($s7) # equilavent C code :- B[2] = h + A[4]

7)

add $t0,$zero,$t1

8)

0010 0010 0001 1000